OpenStudy (anonymous):

A rectangle has its base on the x axis and its upper two vertices on the parabola y = 6 -x^2. What is the largest area the rectangle can have and what are its dimensions?

OpenStudy (campbell_st):

ok... so the parabola looks like  I assumed the rectangle is symmetrical about the y axis and each vertical side is x units from the origin... so the width of the rectangle is 2x the length is f(x) so the area of the rectangle is \[A = 2x \times f(x) = 2x \times 6 - x^2\] so \[A = 12x - 2x^3\] I'd find the derivative and then solve for x